From SMM@avion.stsci.edu Wed Feb 10 13:06:32 1993 Received: Wed, 10 Feb 93 13:06:32 EST from STSCI.EDU by marian.stsci.edu (4.1) Received: Wed, 10 Feb 93 13:06:10 EST from AIRY.STSCI.EDU by stsci.edu (4.1) Received: from DECNET-MAIL (SMM@STOSC) by avion.stsci.edu (PMDF #3144 ) id <01GUK31EHIJKCDR2IZ@avion.stsci.edu>; Wed, 10 Feb 1993 13:04:04 EDT Date: 10 Feb 1993 13:04:04 -0400 (EDT) From: On-Duty SMM x5074 Subject: HST STATUS REPORT 2/09-2/10 To: reppert@stsci.edu, KINNEY@stsci.edu, MDS@stsci.edu, HST-STATUS@stsci.edu Message-Id: <01GUK31EIBHECDR2IZ@avion.stsci.edu> X-Vms-To: @[.DIS]HST_STAT Mime-Version: 1.0 Content-Transfer-Encoding: 7BIT Content-Length: 4232 X-Lines: 112 Status: RO HUBBLE SPACE TELESCOPE DAILY REPORT #820 PERIOD COVERED: 0800 EST 02/09/93 - 0800 EST 02/10/93 Daily Status Report as of 041/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed GHRS/2 2537 (Physical Conditions of Interstellar Gas) The High Resolution Spectrograph was used to make a series of 23 observations in the direction of star HD80009 to obtain high signal to noise ratio spectra to study the interface between the warm and cold regions in the interstellar medium. All of the wave length calibrations and spectral observations were reported to look okay. 1.2 Completed PC 4347 (Flood Retention Test) The Planetary Camera was used to take three Earth flat images which will be used to determine how much of the last ultraviolet flood operation still remain. All three observations were received and displayed. The first image was underexposed, the second was okay, and the last was saturated. 1.3 Completed GHRS/2 3463 (Highly Ionized Nitrogen in Galactic Halo) The High Resolution Spectrograph was used to obtain a total of 11 observations of Class O-7 halo type star HD156359 to obtain absorption line spectra of Nitrogen V. The confirmation images in the large and small science apertures looked good, with the target well centered in the large science aperture and only slightly off center in the small aperture. The wave length calibration and target spectra observations looked good.. 1.4 Completed GHRS/2 1179 (Hybrid Star winds and Transition Regions) The High Resolution Spectrograph was used to make six spectral observations of the star Alpha Traui (aka, First magnitude star Aldeberaun, HD150789) to help learn more about the temperatures, densities, out flow, velocities, and mass loss rates for hybrid stars. The observations were received in the OSS, and will be analyzed later. 1.5 Completed WFC-PC 4109 (Rapid Internal Monitor) Both the Wide Field Camera and the Planetary Camera were used to make internal flat observations to be used to check for new forms of contamination (i.e., measles and daisies). The observations were monitored in the OSS, and no problems were reported. 1.6 Completed PC 2944 (Internal Velocity Distribution in Globular Clusters) The Planetary Camera was used to make observations of the globular cluster NGC-104 (aka, 47 Tucanae) to study its internal velocity dispersion. The observations were executed as scheduled, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 8 Successful : 6 (Failed 039/2119Z and 041/0759Z in CT) Scheduled Reacquisitions : 20 Successful : 20 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 14 Successful : 14 2.3 Significant HSTARs: None 2.4 Operations Note: None 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete PC 2944 (Internal Velocity Distribution In Globular Clusters) 3.2 Complete Three Sets of PC 3603 (Snapshots of Protoplanetary Nebulae) 3.3 Complete HSP/UV1 4795 (Frames/ Observation Reset Fix Test) 3.4 Uplink Updated Ephemeris Tables, Per MOSES 2871 3.5 Complete GHRS/2 3737 (Onset of Chromospheric Activity) 3.6 Complete Two Sets of WFC 2227 (Determination of Extragalactic Distance Scale) 3.7 Complete FOS/BL 4059 (Absolute Photometry Monitor) 3.8 Initiate FOC/96 3487 (Cooling Flows in Distant Clusters) 3.9 Initiate WFC 1306 (GTO Low Latitude Parallel Program) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-